I need to write an equation of the line containing the given point and perpendicular to the given line. (-10,8);7x-5y=2
These are the steps that I have done:
solve for y
7x-5y=2
-5y=-7+2
y=7/5x-2/5
the slope of the given line is 7/5
the slope of the new line is -5/7
next step
y-y1=-5/7(x-x1)
y-8=-5/7(x-x1)
y-8=-5/7(x-(-10))
then I needed to write it in this form= y=mx+b and solve for y
y-8=-5/7(x+10)
I was given the rest of the problem but do not understand how the numbers got to be what they were.
y-8=-5/7x-50/7 How did they get the 50?
y=5/7x+6/7 and the same here, How did they get 6/7?
I have tried to figure this out by still do not understand. Could you please explain how those numbers were reached?

Answer by chessace(471)

"y-8=-5/7(x+10)
I was given the rest of the problem but do not understand how the numbers got to be what they were.
y-8=-5/7x-50/7 How did they get the 50?"
The (-5/7) was distributed over (x+10), the 2nd part of which (+10) results in -5*10/7 = -50/7
"y=5/7x+6/7 and the same here, How did they get 6/7?"
[Note the leading "-" is missing]
8 was added to both sides, 8 = 56/7, - 50/7 = 6/7